I really need some help with this question. Can anyone explain it to me?
Thanks!


The volume V of a square based pyramid with base sides s and height h is:

V = 1/3 s^2 h

If the height is half of the length of a base side, express the volume V as a function of s.

Answer by fractalier(2101)

We are given the formula for the volume
V = 1/3 * s^2 * h
Now this formula gives us the volume in terms of BOTH the side and the height.
If we know the height is equal to half the side, we can write
h = (1/2)s
That allows us to substitute (1/2)s in for h in the original formula so that we have the volume in terms of ONLY the side...thus we get
V = 1/3 * s^2 * h
V = 1/3 * s^2 * (1/2)s
V = (1/6)s^3